arricandy:

Gina has 26 more barrettes than Holly. The equation g = 26 + h, where g represents the number of barrettes Gina has and h represents the number of barrettes Holly has, shows this relationship. If Gina has 49 barrettes, how many barrettes does Holly have?

carmelle:

Since we know Gina has 49 barrettes, we can determine that g = 49 49 = 26 + h Do you know what the next step would be? Are you familiar with solving for variables?

arricandy:

@carmelle wrote:
Since we know Gina has 49 barrettes, we can determine that g = 49 49 = 26 + h Do you know what the next step would be? Are you familiar with solving for variables?
nah g

carmelle:

thats ok, our next step here would be to `subtract`. We are subtracting 26 from both sides so.. 49 - 26 = ? 26 - 26 = ?

arricandy:

@carmelle wrote:
thats ok, our next step here would be to `subtract`. We are subtracting 26 from both sides so.. 49 - 26 = ? 26 - 26 = ?
0

carmelle:

ok.. how about 49 - 26 = ?

CheeseStick7793:

The answer is D. 22 because 47 - 25= 22

arricandy:

@cheesestick7793 wrote:
The answer is D. 22 because 47 - 25= 22
thanks

CheeseStick7793:

or 26

carmelle:

@cheesestick7793 wrote:
The answer is D. 22 because 47 - 25= 22
thats incorrect

arricandy:

@arricandy wrote:
@cheesestick7793 wrote:
The answer is D. 22 because 47 - 25= 22
thanks
Wrong

CheeseStick7793:

i said 26 too

carmelle:

There is only 1 correct answer

CheeseStick7793:

or 23

carmelle:

are you just guessing?

CheeseStick7793:

nope i think it be 23

arricandy:

its right ty

arricandy:

this is my next one

arricandy:

Rewrite the following equation using the distributive property: 4(3 + 5)

carmelle:

49 = 26 + h - 26 -26


26 cancels out because 26 - 26 = 0 to find the value of h, all you have to do is find 49 - 26 :/

arricandy:

@carmelle wrote:
49 = 26 + h - 26 -26 26 cancels out because 26 - 26 = 0 to find the value of h, all you have to do is find 49 - 26 :/
23h

carmelle:

yes, 23 = h

CheeseStick7793:

i said 23

arricandy:

4(3 + 5)

carmelle:

@cheesestick7793 wrote:
i said 23
yea, but we cant give direct answers

CheeseStick7793:

bruh

arricandy:

lmaooooo

arricandy:

carmelle would u date me

carmelle:

distributive property is basically a(b + c) = ab + ac so your problem says 4(3 + 5) what do you think u would do here?

arricandy:

@carmelle wrote:
distributive property is basically a(b + c) = ab + ac so your problem says 4(3 + 5) what do you think u would do here?
add

carmelle:

noo... you multiply.. look the picture I am going to attach and see if that helps

1 attachment
arricandy:

ayyyy yall wanna be friends

carmelle:

do you understand what I am telling you?

arricandy:

@carmelle wrote:
do you understand what I am telling you?
nah

carmelle:

what dont you understand so that I can help? :)

CheeseStick7793:

wht

CheeseStick7793:

is

CheeseStick7793:

going

CheeseStick7793:

on

CheeseStick7793:

??

arricandy:

@carmelle wrote:
what dont you understand so that I can help? :)
i want chciken

carmelle:

... 4(3 + 5) When distributing, you are multiplying whats outside the parenthesis with what's inside. 4(3) + 4(5) 4 × 3 = ? 4 × 5 = ? Once you get your answers, you can add the numbers together and you will get your final answer.

arricandy:

daddy